OpenStudy (anonymous):

Help Needed : (2log6 + 6log2) / (4log2 + log 27 - log 9 )

OpenStudy (cwrw238):

apply laws of logs log a + log b = log ab log a - log b = log a/b log a^n = n log a

OpenStudy (cwrw238):

so applying this to the numerator 2 log 6 = log 6^2 = log 36 6 log 2 = log 2^6 = log 64 log 36 + log 64 = log 36*64

OpenStudy (cwrw238):

can you try simplifying the denominator?

OpenStudy (cwrw238):

what about 4 log 2?

OpenStudy (anonymous):

i get : log 2304 base 48 .. :/

OpenStudy (cwrw238):

this is the third law in reverse n log a = log a^n for 4 log 2 n = 4 and a = 2

OpenStudy (anonymous):

Got the answer ! :D Thankkyooou :D

OpenStudy (cwrw238):

log 36*64 = log 2304 - right

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Ans = 2 :)